Consider the following substances: CH5N, C2H5O2COOH, H3AsO4, KClO2, HC2H2ClO2. For each compound Decide whether the co...

  1. Consider the following substances: CH5N, C2H5O2COOH, H3AsO4, KClO2, HC2H2ClO2. For each compound
    1. Decide whether the compound behaves as an acid or base in solution
    2. Write the equation for the dissociation of the compound in water
    3. Write the equilibrium expression for the dissociation in water

  1. Calculate the [H+], the [OH-], the pH and the pOH for a solution prepared by diluting 50.0 ml of 0.035 M HNO3 to 250 ml in a volumetric flask.

  1. Fluoroacetic acid, C2H3FO2, occurs naturally in at least 40 plants found in Australia, Brazil and in Africa, notably in “Gifblaar” which is one of the most poisonous plants found in South Africa. The compound is highly toxic and causes convulsions and ventricular fibrillation and has been used to poison rats. If the Ka of this acid is 2.57 x 10-3 determine the [H+], the [OH-], the pH and the pOH for a 0.05 M solution of this acid.
Answer #1

Answer:

a, b CH5N + H2O → CH5NH+ + OH- Act as base

C2H5O2COOH + H2O → C2H5O2COO- + H3O+ Act as acid

H3AsO4 + H2O → H2AsO4- + H3O+ Act as acid

KClO2 + H2O → HClO2 + KOH It's a salt. But HClO2 is weak acid and KOH is strong base. The solution is little basic.

ClCH2COOH + H2O → ClCH2COO- + H3O+ Act as acid

c. The initial concentration of HNO3 = 0.035 M

Final Volume = 250 mL

The final concentrations is = (0.035 x 50)/250 = 7 x 10-3 M

As HNO3 is strong acid, the dissociation is ~100 %. So, the concentration of [H+] = 7 x 10-3 M

Hence, pH = -log( 7 x 10-3) = 2.15

and pOH 14 - 2.15 = 11.85

[OH-] = 10-11.85 = 1.41 x 10-12 M

d. Here, HA is an acid = Fluoroacetic acid [C2H3FO2]

HA --> H+ + A-
Initial 0.05 0 0
Change -a +a +a
Equilibrium 0.05 - a a a

Ka = [H+][A-]/[HA] = 2.57 x 10-3

2.57 x 10-3 = a.a/(0.05 -a)

or a2 + 2.57 x 10-3a - 1.285 x 10-4 = 0

[H+] = 0.010123 M

pH = 1.99

pOH = 12.01 and [OH-] = 9.77 x 10-13 M

Thanks